Steam flows steadily through an adiabatic turbine. The inlet conditions of the steam are 10 MPa,
450
1
Expert's answer
2021-09-30T15:28:27-0400

"P_1= 10 \;MPa \\

T_1 = 450 \;°C \\

h_1 = 3240.9 \;kJ/kg"

At "P_2=10 \;kPa" and x=0.92, from table

"h_f = 191.83 \;kJ/kg \\

h_{fg} = 2392.8 \;kJ/kg \\

h_2 = h_p + h_{fg} = 191.83 +0.92(2392.8) = 2393.206 \;kJ/kg \\

Q-W = m[(h_2-h_1) + \frac{v^2_2 -v^2_1}{2000}] \\

Q=0 \\

W = m[(h_1-h_2) + \frac{v^2_1 -v^2_2}{2000}] \\

= 14 [(3240.9 -2393.206) + \frac{85^2 -54^2}{2000}] \\

= 11897.88 \;kJ/s \\

= 11.90 \;MW"